Refer to this information. Calculate the future value of a 4 year investment of R13270 at an interest rate of 6.09% per annum compounded quarterly. 

If the interest rate changes at the end of the second year from 6.09$ per annum compounded quarterly to 7.54% per annum compounded quarterly, the accumulated value of the original investment of R13270 at the end of 4 years will be.. 


1
Expert's answer
2021-12-01T18:37:01-0500

"FV = PV(1 + r\/m)^{mt}"


1.

"FV = 13270(1 + 0.0609/4)^{4\cdot4}=R\ 16899.26"


2.

at the end of the second year:

"FV = 13270(1 + 0.0609/4)^{2\cdot4}=R\ 14975.09"

at the end of 4 years:

"FV = 14975.09(1 + 0.0754/4)^{2\cdot4}=R\ 17388.07"